# Designing Simple Mechanisms

### v1

@@ 1 Thinking contingently about unobserved moves
The assumptions we make about preferences affect the structure of OSP mechanisms. Recent work has explored beyond auctions, characterizing OSP mechanisms for settings such as two-sided matching, social choice, and object allocation without transfers (Arribillaga et al., 2020, 2023; Ashlagi and Gonczarowski, 2018; Bade, 2019; Bade and Gonczarowski, 2017; Mandal and Roy, 2022; Thomas, 2020; Troyan, 2019). Other work has built tools to study OSP mechanisms, providing a ‘revelation principle’ (Mackenzie, 2020) and algorithms to construct them (Golowich and Li, 2021).
@@ 2 Planning for future moves
Suppose that we have a set of indivisible goods and a set of agents, and we want to allocate one good to each agent, without money transfers. One natural solution is to order the agents randomly, give the first agent their favorite good, then give the second agent their favorite good among those that remain, and so on iteratively. This algorithm results in an ex post efficient allocation (Bogomolnaia and Moulin, 2001).
We can ask agents to simultaneously submit rank-order lists of the goods, and then process those lists according to the above algorithm. The resulting mechanism—the static random priority mechanism (static RP)—is strategy-proof. However, it is not obviously strategy-proof. If you rank goods truthfully, you might receive your third-favorite, whereas if you deviate to rank your second-favorite good at the top, then you might receive your second-favorite good, which you strictly prefer.
@@
Alternatively, we could approach agents one-by-one in random order, asking each to pick their favorite good among those that remain. The resulting dynamic random priority mechanism (dynamic RP) is obviously strategy-proof.
@@
In practice, dynamic RP results in higher rates of truthful play than its static equivalent. In an incentivized laboratory experiment with four players and four goods, dynamic RP resulted in the dominant-strategy outcome in 93% of games, whereas static RP resulted in the dominant-strategy outcome in only 64% of games (Li, 2017, p. 3282). [^13] Dreyfuss et al. (2022) and Meisner and von Wangenheim (2023) argue that this disparity might be due to loss aversion rather than strategic mistakes.
@@
Ascending auctions and dynamic RP are both intuitively simple. But some games with obviously dominant strategies are not intuitively simple. For example, take any chess position such that White can force a win, and consider the subgame that starts from that position. [^14] It is not known whether White can force a win from the opening position. World Chess Champion Bobby Fischer opined, “I think it’s almost definite that the game is a draw theoretically”. If White plays the win-forcing strategy, then by definition the worst-case outcome is that White wins. If White ever deviates from that strategy, the best-case outcome is that White wins. Thus, the win-forcing strategy is obviously dominant in the subgame. However, real chess players often fail to play the win-forcing strategy, even from positions that appear in an endgame tablebase (Anderson et al., 2017).
@@
What makes dynamic RP simple, and chess complicated? Intuitively, to identify a win-forcing strategy in chess, one has to plan far in advance, consider different future contingencies, and then backward induct in order to judge the merits of each current move. By contrast, when playing dynamic RP, each player chooses exactly once, so forward planning is not required. Conventional game theory elides this distinction, because a strategy is a complete contingent plan, specifying not only what one does right now, but also what one will do at all future contingencies.
@@
Pycia and Troyan (2023) formalized the idea of limited forward planning, as follows: We consider games of perfect recall, that is, the information sets are such that each participant ‘remembers’ all the past information sets that they encountered and all the past moves they took. Let $\mathcal{I}_{n}$ denote the family of information sets assigned to player $n$. Suppose that at each information set $I\in\mathcal{I}_{n}$, player $n$ can only foresee some of the other information sets. We use $F(I)\subseteq\mathcal{I}_{n}$ to denote the family of information sets foreseeable at $I$, and assume that $I\in F(I)$. A partial strategic plan for information set $I$, denoted $S_{n}^{I}$, chooses actions at all information sets in $F(I)$. The plan $S_{n}^{I}$ is $F$-dominant if, upon reaching information set $I$, the worst-case outcome from following $S_{n}^{I}$ is at least as good as the best-case from deviating to another plan that chooses differently at $I$. Specifically, the worst-case and the best-case are with respect to the actions of other players and also the actions that player $n$ takes at unforeseeable information sets, those in $\mathcal{I}_{n}\setminus F(I)$.
@@
This approach results in a variety of incentive criteria, that depend on how we specify the foreseeable information sets $F(I)$. At one extreme, all information sets are foreseeable, that is, $F(I)=\mathcal{I}_{n}$, and $F$-dominance is equivalent to obvious dominance. On the other extreme, only the present information set is foreseeable, so $F(I)=\{I\}$, and we could call the resulting criterion strong obvious dominance. For example, in dynamic RP, picking your favorite object is strongly obviously dominant. Each player is called to play exactly once, so it does not matter that future information sets are unforeseeable.
@@
Sometimes optimal behavior requires an intermediate degree of foresight. Consider a bidder who values the object at \$40. They are participating in an ascending auction, with the price rising by \$1 in each step. The price starts at \$1, and the bidder reasons, “If I quit now my payoff is \$0 for sure. But if I agree to bid \$1 and plan to quit at \$2, then my payoff could be \$$\$40-\$1=\$39$, and is at least \$0. So I’ll keep going for now.” Then the price rises to \$2, and the bidder thinks, “If I quit now my payoff is \$0 for sure. But if I keep bidding, planning to quit at \$3, then my payoff could be \$$\$40-\$2=\$38$, and is at least \$0.” Thus, by looking one step ahead at each point, the bidder is led to a pattern of behavior that reproduces the optimal strategy. Notice that the bidder is making a collection of partial strategic plans, and these plans are not consistent with each other. When the price is \$1, they plan to quit at \$2. But when the price reaches \$2, they revise that plan.
@@
To formalize this idea, let the foreseeable information sets $F(I)$ consist of $I$ and every information set that is ‘one step ahead’ of $I$, and let us call the resulting incentive criterion one-step dominance. [^15] That is, where the same player is called to play, without any other such information sets in between. Observe that the ascending auction is one-step simple, meaning that truthful bidding can be induced by a collection of partial strategic plans (one for each information set), each of which is one-step dominant. Just a little bit of foresight is enough to play optimally in an ascending auction.
@@
Let us return to the setting with indivisible goods and no transfers. As we saw, dynamic RP is OSP. But there are other OSP mechanisms in this setting; Pycia and Troyan (2023) characterized those mechanisms, and some of them are not intuitively simple, instead requiring participants to plan far into the future in order to see that the mechanism is incentive-compatible. They showed that strengthening the requirement from OSP to one-step simplicity rules out the counter-intuitive mechanisms.
@@
It can take detailed contingent thinking to see that a revelation mechanism, such as the second-price auction, is incentive-compatible. Dynamic mechanisms, such as the ascending auction, can mitigate this difficulty by paring down the contingencies that participants need to consider. [^16] Glazer and Rubinstein (1996) formalize another sense in which dynamic games can be easier to understand than their static counterparts. But dynamic mechanisms can raise new difficulties, requiring participants to plan far in advance in order to understand that a given strategy is optimal. As we saw, the concept of partial strategic plans introduced by Pycia and Troyan (2023) provides a way to study this kind of complexity.
@@ 3 Reasoning about other players’ beliefs
The study of simplicity is not only about refining strategy-proofness. Even mechanisms without dominant strategies can vary in how simple they are.
In some mechanisms, each player has to reason about the other players’ beliefs. Consider bilateral trade with transfers; there is a seller who can produce an indivisible object, and a prospective buyer. The seller has a privately-known cost $C$ and the buyer has a privately-known value $V$. If a sale is made at some price $p$, then the buyer’s payoff is $V-p$ and the seller’s payoff is $p-C$. Otherwise both payoffs are 00.
One natural mechanism for bilateral trade is the double auction, studied by Chatterjee and Samuelson (1983): The seller and buyer simultaneously submit offers, $s$ and $b$ respectively. They transact if and only if $s\leq b$, at price $\alpha s+(1-\alpha)b$, for parameter $\alpha\in[0,1]$. Suppose for the moment that $\alpha=0.5$, so the mechanism splits the difference between the two offers. Then neither side has a dominant strategy. The seller’s utility-maximizing offer depends on the buyer’s offer $b$; if the seller knew $b$, then they would set $s=b$ if $C\leq b$, and set $s>b$ otherwise.
@@
This reasoning process seems fantastical, and in practice will be beyond the capacity of many participants. [^17] In many theoretical analyses, we abstract away from this subtlety by studying Bayesian Nash equilibria and making strong assumptions about higher-order beliefs; for instance, that both buyer and seller have a common prior about the distribution of values. These simplifying assumptions are often useful. There is a recent literature that seeks to relax them (Bergemann and Morris, 2005; Bergemann et al., 2013). If participants are inexperienced, it seems unlikely that they will find the equilibrium of the double auction with parameter $\alpha=0.5$ from first principles.
@@
Some mechanisms do not require participants to reason about higher-order beliefs. Suppose that we take the double auction, and instead set the parameter $\alpha$ equal to $1$. Then the seller is effectively making a take-it-or-leave-it offer of $s$ to the buyer, and the transaction occurs at a price of $s$ if and only if $s\leq b$. It is a dominant strategy for the buyer to set $b=V$, so the buyer does not need to think about the seller’s strategy. If the seller knows that the buyer will set $b=V$, then the seller’s problem reduces to choosing an offer $s$ to maximize $(s-C)P(s\leq V\mid C)$. Observe that the solution depends on the seller’s cost $C$ and on the seller’s belief about the buyer’s value $V$, but not on any higher-order beliefs. The seller can calculate their optimal offer using only their first-order beliefs, although they do not have a dominant strategy.
@@
Börgers and Li (2019) proposed a new incentive criterion, picking out mechanisms in which participants can ‘play well’ using just their first-order beliefs. The proposal is as follows: Let us fix, for each player $n$, a set of possible utility functions $\mathcal{U}_{n}$ (e.g. the different values that the buyer might have for the object). A strategy for player $n$ is a function from $\mathcal{U}_{n}$ to distributions over actions (e.g. in the double auction, the actions are the feasible offers). A strategy is undominated if there is no other strategy that is always at least as good, and sometimes strictly better. A first-order belief $\mu_{n}$ for player $n$ is a distribution over the other players’ utility functions, that is $\mu_{n}\in\Delta(\mathcal{U}_{-n})$. For instance, the seller might believe that the buyer’s value is uniformly distributed on $\{0,1,2,\ldots,10\}$. A strategy $S_{n}$ is robust with respect to first-order belief $\mu_{n}$ if, assuming the other players’ utilities are distributed according to $\mu_{n}$, the strategy $S_{n}$ is a best response to every profile of undominated strategies for the other players. A mechanism is strategically simple if for every player $n$ and every first-order belief $\mu_{n}\in\Delta(\mathcal{U}_{-n})$, there exists a robust strategy with respect to $\mu_{n}$.
@@
Every strategy-proof mechanism is strategically simple, but not vice versa. For example, the double auction with $\alpha=1$ is strategically simple. For the buyer, offering $b=V$ is a dominant strategy, and hence a robust strategy. Moreover, offering $b=V$ is the buyer’s only undominated strategy, so the seller has a robust strategy that sets $s$ to maximize $(s-C)P(s\leq V\mid C)$, which depends only on the seller’s cost $C$ and on the seller’s beliefs about the distribution of $V$, i.e. the seller’s first-order belief. By contrast, the double auction with $\alpha=0.5$ is not strategically simple. The buyer has many undominated strategies, and for non-trivial seller beliefs there is no robust seller strategy.
@@
Börgers and Li (2019) proved that in the bilateral trade setting, all the strategically simple mechanisms have a special structure: The ‘leading’ player either chooses an offer $p$ from a set of permitted prices, or declines trade. If an offer of $p$ is made, then the ‘following’ player either accepts (and trade occurs at price $p$) or declines, in which case no trade occurs. This class includes the double auction with $\alpha=0$ and with $\alpha=1$, as well as variations that restrict the set of permitted prices.
@@
The class of local dictatorships is subtle, and includes mechanisms that are not ‘dictatorial’ in the colloquial sense. One example of a local dictatorship is a voting mechanism in which each committee member submits a list ranking all alternatives and (at the same time) the chairperson selects two alternatives for a run-off vote. From the selected pair, the final outcome is the alternative that is ranked higher by a majority of the committee. (Suppose that all players have strict preferences over the alternatives.) Fixing a profile of utility functions, each committee member has only a single undominated action, namely, to submit their truthful rank-order list. Thus, in the restricted mechanism only the chairperson’s action determines the outcome.

### v3

@@ 2 Thinking contingently about unobserved moves by other players
The assumptions we make about preferences affect the structure of obviously strategy-proof mechanisms. Recent work has explored beyond auctions, characterizing obviously strategy-proof mechanisms for settings such as two-sided matching, social choice, and object allocation without transfers (Arribillaga et al., 2020, 2023; Ashlagi and Gonczarowski, 2018; Bade, 2019; Bade and Gonczarowski, 2017; Mandal and Roy, 2022; Thomas, 2020; Troyan, 2019). Other work has built tools to study obviously strategy-proof mechanisms, providing a ‘revelation principle’ (Mackenzie, 2020) and algorithms to construct them (Golowich and Li, 2021).
@@ 3 Planning for future moves
Some mechanisms require participants to plan far in advance for their own future moves, in order to see that the mechanism is incentive-compatible.
@@
Suppose that we have a set of indivisible goods and a set of agents, and we want to allocate one good to each agent, without money transfers. One natural solution is to order the agents randomly, give the first agent their favorite good, then give the second agent their favorite good among those that remain, and so on iteratively. This algorithm results in an ex post efficient allocation: any other allocation that makes one agent better off, must also make another agent worse off (Bogomolnaia and Moulin, 2001).
@@
To carry out the above algorithm, we need to ask agents about their preferences. We could approach agents one-by-one in random order, asking each to pick their favorite good among those that remain. The resulting dynamic random priority mechanism is obviously strategy-proof.
In contrast, we could ask each agent to submit rank-order lists of the goods, and then process those lists according to the algorithm. The resulting static random priority mechanism is strategy-proof. However, it is not obviously strategy-proof. Because of the randomness of the order in which the rank-order lists are chosen, it is possible that if you rank goods truthfully, you might receive your third-favorite, whereas if you deviate to rank your second-favorite good at the top, then you might receive your second-favorite good, which you strictly prefer.
@@
In practice, dynamic random priority results in higher rates of truthful play than its static equivalent. In an incentivized laboratory experiment with four players and four goods, dynamic random priority resulted in the dominant-strategy outcome in 93 percent of games, whereas static random priority resulted in the dominant-strategy outcome in only 64 percent of games (Li, 2017, p. 3282). [^8] Dreyfuss et al. (2022) and Meisner and von Wangenheim (2023) argue that this disparity might be due to loss aversion rather than strategic mistakes.
@@
Ascending auctions and dynamic random priority are both intuitively simple. But some games with obviously dominant strategies are not intuitively simple. For example, take any chess position such that White can force a win, and consider the subgame that starts from that position. If White plays the win-forcing strategy, then by definition the worst-case outcome is that White wins. If White ever deviates from that strategy, then the best-case outcome is that White wins. Thus, the win-forcing strategy is obviously dominant in the subgame. However, real chess players often fail to play the win-forcing strategy, even from positions where such a strategy has been found by computers (Anderson et al., 2017). (One of the unanswered questions of chess is whether white can force a win from the opening position. Chess great Bobby Fischer opined, “I think it’s almost definite that the game is a draw theoretically.”)
@@
What makes dynamic random priority simple, and chess complicated? Intuitively, to identify a win-forcing strategy in chess, one has to plan far in advance, consider different future contingencies, and then backward-induct to judge the merits of each current move. By contrast, when playing dynamic random priority, each player chooses exactly once, so forward planning is not required. Conventional game theory elides this distinction, because a strategy is viewed as a complete contingent plan, specifying not only what one does right now, but also what one will do at all future contingencies.
@@
One way that a mechanism can be simple is that it might require only limited forward planning. Pycia and Troyan (2023) formalized the idea of limited forward planning. They consider games of perfect recall, that is, the information sets are such that each participant “remembers” all the past information sets that they encountered and all the past moves they took. However, at each information set, the active player can only foresee some of their other information sets. Thus, each player forms a partial strategic plan based on their own information set, that specifies moves at foreseeable information sets. This plan is simply dominant if the worst-case outcome from following the plan is at least as good as the best-case from deviating to another plan that chooses differently at that information set. Specifically, the worst-case and the best-case are with respect to the actions of other players and also the actions that the player takes at unforeseeable information sets.
@@
This approach to modeling limited forward planning results in a variety of incentive criteria, which depend on how we specify the foreseeable information sets. At one extreme, suppose that all information sets are foreseeable; in this case, simple dominance is equivalent to obvious dominance. At the other extreme, suppose that only the present information set is foreseeable, and let us call the resulting criterion strong obvious dominance. For example, in dynamic random priority, picking your favorite object is strongly obviously dominant. Each player is called to play exactly once, so it does not matter that future information sets are unforeseeable. Strong obvious dominance may seem too demanding a criterion, but there is a large literature studying the welfare and revenue guarantees of posted-price mechanisms, which have strongly obviously dominant strategies (Lucier, 2017).
@@
However, one can imagine an intermediate case, where the player can foresee some, but not all, of the future information sets. Consider a bidder who values the object at \$40. The bidder is participating in an ascending auction, with the price rising by \$1 in each step. The price starts at \$1, and the bidder reasons, “If I quit now my payoff is \$0 for sure. But if I agree to bid \$1 and plan to quit at \$2, then my payoff could be \$$\$40-\$1=\$39$, and is at least \$0. So I’ll keep going for now.” Then the price rises to \$2, and the bidder thinks, “If I quit now my payoff is \$0 for sure. But if I keep bidding, planning to quit at \$3, then my payoff could be \$$\$40-\$2=\$38$, and is at least \$0.” Thus, by looking just one step ahead at each point, the bidder is led to behave in a way that reproduces the optimal strategy. Notice that the bidder is making a collection of partial strategic plans, and these plans are not consistent with each other. When the price is \$1, they plan to quit at \$2. But when the price reaches \$2, they revise that plan.
@@
To formalize this idea, let the foreseeable information include both the current information set (which includes all the player’s earlier moves) and every information set that is ‘one step ahead.’ Observe that the ascending auction is one-step simple, meaning that truthful bidding can be induced by a collection of partial strategic plans (one for each information set), each of which is simply dominant when the player looks just one step ahead. Just a little bit of foresight is enough to play optimally in an ascending auction.
@@
Let us return to the allocation of indivisible goods without transfers. As we saw, dynamic random planning is obviously strategy-proof. But there are other obviously strategy-proof mechanisms in this setting; Pycia and Troyan (2023) characterized those mechanisms, and some of them are not intuitively simple, instead requiring participants to plan far into the future in order to see that the mechanism is incentive-compatible. They showed that strengthening the simplicity requirement from obvious strategy-proofness to one-step simplicity rules out the counterintuitive mechanisms.
@@
It can take detailed contingent thinking to see that a static mechanism, such as the second-price sealed-bid auction, is incentive-compatible. Dynamic mechanisms, such as the ascending auction, can mitigate this difficulty by paring down the contingencies that participants must consider. But dynamic mechanisms can raise new difficulties, in some cases requiring participants to plan in advance—even far in advance—to understand that a given strategy is optimal. The concept of partial strategic plans formally introduced by Pycia and Troyan (2023) provides a way to study this kind of complexity. [^9] Jéhiel (1995) and Jéhiel and Samet (2007) also study limited forward planning in games, but it is an open question whether one can adapt those ideas for use in mechanism design.
@@ 4 Reasoning about other players’ beliefs
The study of simplicity is not only about refining strategy-proofness. Even mechanisms without dominant strategies can vary in how simple they are. In particular, some mechanisms require participants to reason about other players’ beliefs, in order to see that the mechanism is incentive-compatible.
Consider bilateral trade with transfers; there is a seller who can produce an indivisible object, and a prospective buyer. The seller has a privately-known cost $C$ and the buyer has a privately-known value $V$. If a sale is made at some price $p$, then the buyer’s payoff is $V-p$ and the seller’s payoff is $p-C$. Otherwise, both payoffs are 00.
One natural mechanism for bilateral trade is the double auction, studied by Chatterjee and Samuelson (1983): The seller and buyer simultaneously submit offers, $s$ and $b$ respectively. They transact if and only if the offer $s$ from the seller is greater than or equal to the offer $b$ from the buyer. The resulting price is a weighted average of $s$ and $b$, that is, $\alpha s+(1-\alpha)b$ for parameter $\alpha\in[0,1]$. Suppose for the moment that $\alpha=0.5$, so the mechanism splits the difference between the two offers. In this case, neither side has a dominant strategy. The seller’s utility-maximizing offer depends on the buyer’s offer $b$; if the seller knew the buyer’s offer $b$, then the seller would set $s=b$ if the seller’s cost satisfies $C\leq b$, and set $s>b$ otherwise.
@@
This many-level reasoning process seems fantastical, and in practice will be beyond the capacity of many participants. If participants are inexperienced, it seems unlikely that they will find the equilibrium of the double auction with parameter $\alpha=0.5$ from first principles.
@@
However, some mechanisms do not require participants to reason about higher-order beliefs. Suppose that we take the double auction, and instead set the split-the-difference parameter $\alpha$ equal to $1$. Then the seller is effectively making a take-it-or-leave-it offer of $s$ to the buyer, and the transaction occurs at a price of $s$ if and only if $s\leq b$. In this case, it is a dominant strategy for the buyer to set $b=V$, so the buyer does not need to think about the seller’s strategy. If the seller knows that the buyer will set $b=V$, then the seller’s problem reduces to choosing an offer $s$ to maximize $(s-C)$ multiplied by the probability that $s\leq V$. Observe that the solution depends on the seller’s cost $C$ and on the seller’s belief about the buyer’s value $V$, but not on any higher-order beliefs. The seller can calculate an optimal offer using only first-order beliefs, although the seller does not have a dominant strategy.
@@
Börgers and Li (2019) proposed a new incentive criterion, picking out mechanisms in which participants can ‘play well’ using just their first-order beliefs. The proposal is as follows: Let us fix, for each player, a set of possible utility functions (which can be thought of as the different values that each buyer might have for an object). Each player will then form a strategy, which means choosing an action based on their utility function: for example, in the double auction, the actions are the feasible offers. A strategy is undominated if there is no other strategy that is always at least as good, and sometimes strictly better. A first-order belief is a distribution over the utility functions of the other players: for instance, in a double auction, the seller might have a first-order belief that the buyer’s value is uniformly distributed on all integers between 0 and 10. A strategy is robust with respect to that first-order belief if, assuming the other players’ utilities are indeed distributed according to that belief, the strategy is a best response to every profile of undominated strategies for the other players. A mechanism is strategically simple if for every player and every first-order belief, there exists a robust strategy.
@@
Every strategy-proof mechanism is strategically simple, but the reverse does not hold true. For example, the double auction with split-the-difference parameter $\alpha=1$ is strategically simple but not strategy-proof. For the buyer, offering $b=V$ is a dominant strategy, and hence a robust strategy. Moreover, offering $b=V$ is the buyer’s only undominated strategy, so the seller has a robust strategy which depends only on the seller’s cost $C$ and on the seller’s beliefs about the distribution of $V$, i.e. the seller’s first-order belief. By contrast, the double auction with $\alpha=0.5$ is not strategically simple. The buyer has many undominated strategies, and for non-trivial seller beliefs there is no robust seller strategy.
@@
Börgers and Li (2019) formally proved that in the bilateral trade setting, all the strategically simple mechanisms have a special structure: The ‘leading’ player either chooses an offer $p$ from a set of permitted prices, or declines trade. If an offer of $p$ is made, then the ‘following’ player either accepts (and trade occurs at price $p$) or declines, in which case no trade occurs. This class includes the double auction with $\alpha=0$ and with $\alpha=1$, as well as variations that restrict the set of permitted prices.
@@
The meaning of what they call “local dictatorships” is subtle, and includes mechanisms that are not “dictatorial” in the colloquial sense. One example of a local dictatorship is a voting mechanism in which each committee member submits a list ranking all alternatives and (at the same time) the chairperson selects two alternatives for a head-to-head vote. From the selected pair of outcomes, the final outcome is the alternative that is ranked higher by a majority of the committee (suppose that all players have strict preferences over the alternatives). In this situation, each committee member has only a single undominated action, namely, to submit their truthful rank-order list. Thus, in the restricted mechanism, the chairperson’s action determines the outcome.